A certain scale only registers weights that are greater than 6 pounds. A person who wanted to know the weights of a puppy, a kitten, and a bunny weighted them in pairs and go the following results.
The kitten and the bunny weighted 7 pounds.
The kitten and the puppy weighted 8 pounds.
The bubby and the puppy weighed 9 pounds.
What is the weight of the puppy?
A. 2 pounds B. 3 pounds C. 4 pounds D. 5 pounds E. 6 pounds

Answer by stanbon(75887)

The kitten and the bunny weighted 7 pounds.
The kitten and the puppy weighted 8 pounds.
The bubby and the puppy weighed 9 pounds.
What is the weight of the puppy?
--------------
Equations:
k + b = 7
k + p = 8
b + p = 9
-----------------
Rearrange:
K + b +0 = 7
k + 0 +p = 8
0 + b +p = 9
----------------
Solve with any method you know to get:
k = 3
b = 4
p = 5
===========
Cheers,
Stan H.
===========
A. 2 pounds B. 3 pounds C. 4 pounds D. 5 pounds E. 6 pounds